Identify the y intercept,the vertex,and graph:

y=2(x-3)squared-4
I cannot figure out how to make the #2 squared on the computer keys!hopefully this understandable

Answer by rfer(16322)

y=2(x-3)^2-4
y=2(x-3)(x-3)-4
y=2(x^2-6x+9)-4
y=2x^2-12x+18-4
y=2x^2-12x+14